If line segment JK is parallel to the line segment LM, ∠JKM = 2x²+15 and ∠LMK=29x-45, find ∠JKM. There is a parallelogram KLMJ with diagonal KM.
1
Expert's answer
2010-07-07T05:54:19-0400
2x²+15 = 29х -45 => x1 = 5/2; x2 = 12 => ∠JKM = 2*144+15 = 303 - it is impossible, because angle is sharp, or ∠JKM = 2*25/4 + 15 = 55/2 - this is an answer.